a :Ñ:N: b. :C:C:::N: H:C::F 30: :0: l H:ö::[::ö: HExamine each of the following electron-dot formulas and decide whether the formula is correct, or whether you could write a formula that better approximates the electron structure of the molecules. State which concepts or rules you use in each case to arrive at your conclusion.
Answer #1

ANS-:formal charge: The charge assigned to an atom in a molecule, assuming that electrons in a chemical bond are shared equally between atoms. This helps determine which of a few Lewis structures is most correct.

Determining Formal Charge-:

*Although we know how many valence electrons are present in a compound, it is harder to determine around which atoms the electrons actually reside. To assist with this problem, chemists often calculate the formal charge of each atom. The formal charge is the electric charge an atom would have if all the electrons were shared equally.

*The formal charge of an atom can be determined by the following formula:

************[FC=VE−(NE+BE/2)]

*********In this formula, VE represents the number of valence electrons of the atom in isolation, NE is the number of non-bonding valence electrons, and BE is the total number of electrons in covalent bonds with other atoms in the molecule.

3222 (0) I left Nihogen tomul charge = VE - NE - M: M: Comect Lewis formulas : and better formula le formal charge of each NiVE em outermost electron in Oxyge formal charge VE=6, NE=8 BE=0 fomal charge = 6-8-0 = -2 Now Correct formula r. 72- 6 - formLESS FORMAL CHARGE ,HIGHER AND BETTER ELECTRON DOT FORMULA OR LEWIS DOT STRUCTURE THANK YOU VERY MUCH.GOOD LUCK.